# Randomized Physics-Informed Machine Learning for Uncertainty Quantification in High-Dimensional Inverse Problems

### 11 Dec 2023 v1

@@ Abstract
We propose a physics-informed machine learning method for uncertainty quantification (UQ) in high-dimensional inverse problems. In this method, the states and parameters of partial differential equations (PDEs) are approximated with truncated conditional Karhunen-Loève expansions (CKLEs), which, by construction, match the measurements of the respective variables. The maximum a posteriori (MAP) solution of the inverse problem is formulated as a minimization problem over CKLE coefficients where the loss function is the sum of the norm of PDE residuals and $\ell_{2}$ regularization term. This MAP formulation is known as the physics-informed CKLE (PICKLE) method. Uncertainty in the inverse solution is quantified in terms of the posterior distribution of CKLE coefficients, and we sample the posterior by solving a randomized PICKLE minimization problem, formulated by adding zero-mean Gaussian perturbations in the PICKLE loss function. We call the proposed approach the randomized PICKLE (rPICKLE) method. The noise variance of perturbations in the regularization term is given by the prior (standard normal) distribution of CKLE coefficients. The noise variance in the residual term, $\sigma^{2}_{r}$, becomes the only free parameter in the rPICKLE method, whose value is determined to maximize the log predictive probability (LPP) of the parameter predictions.
We test rPICKLE for the inverse problems of estimating parameters and states in groundwater models described by the diffusion (Darcy) equation with low and high-dimensional parameter space. We validate rPICKLE for the low-dimensional case with 15 unknown CKLE parameters by showing that rPICKLE and Hamiltonian Monte Carlo (HMC) produce similar posterior distributions. However, rPICKLE execution time is practically independent of the residual noise variance $\sigma^{2}_{r}$ and is orders of magnitude faster than that of HMC, whose cost also increases with decreasing $\sigma^{2}_{r}$. The execution times of both methods increase with the dimensionality of the problem. However, the execution time of HMC increases significantly faster with the problem dimensionality (and decreasing $\sigma^{2}_{r}$) than that of rPICKLE. For the high-dimensional case (2000 CKLE parameters) with small $\sigma^{2}_{r}$, HMC does not reach the stopping criterion (the set number of samples) after running the code for 30 days. On the other hand, rPICKLE generates the same number of samples in four to five days depending on the value of $\sigma^{2}_{r}$. The posterior mean agrees well with the MAP prediction given by PICKLE for a wide range of $\sigma^{2}_{r}$, and the reference solution lies mostly within the 95% credible interval computed from rPICKLE, indicating that rPICKLE posterior is highly informative.
We also demonstrate that increasing dimensionality and decreasing $\sigma^{2}_{r}$ leads to increasing condition number of the posterior covariance matrix, which is known to decrease the efficiency of HMC. On the other hand, rPICKLE is found to be robust with respect to the condition number and, therefore, is well suited for high-dimensional problems with strict physics constraints, i.e., small $\sigma^{2}_{r}$.
@@ 1 Introduction
Inverse uncertainty quantification (IUQ) problems are ubiquitous in the modeling of natural and engineering systems governed by partial differential equations (PDEs) (e.g., subsurface flow systems, geothermal systems, CO2 sequestration, climate modeling). IUQ has the same challenges as forward UQ, including the curse of dimensionality (COD), and additional challenges associated with the non-uniqueness of inverse PDE problems Zhou et al. (2014); Linde et al. (2017).
@@
In the inverse setting, given the measurements of $y$ and $u$, deterministic inverse solutions give the point estimates of $y$ and $u$. The standard method for finding the point estimate is to solve the PDE-constrained optimization problem where the objective function is the sum of the square differences between predicted and observed $y$ and $u$ values. These problems are ill-posed and a unique solution cannot be found without adding regularity constraints. Adding regularization, which can be interpreted as prior knowledge, leads to the maximum a posteriori (MAP) formulation of the inverse solution. However, the PDE-constrained MAP formulations may suffer from the COD. For example, for the considered in this work diffusion equation model with the unknown space-dependent diffusion coefficient, MAP’s computational cost increases as $O(n^{3})$, where $n$ is the number of unknown parameters (discrete values of the diffusion coefficient) Yeung et al. (2022).
Models of natural and engineering systems can have thousands of (unknown) parameters, and many dimension reduction methods have been proposed for dealing with model complexity Anderson et al. (2015). For unknown parameter fields, the dimension reduction methods usually exploit the spatial correlation of the parameters. For example, in the pilot point method RamaRao et al. (1995); Doherty et al. (2010), parameters are estimated in a few pre-selected locations (pilot points), and the rest of the field is reconstructed using the Gaussian process regression (GPR) or “Kriging” method. Another approach is to linearly project parameters onto a low-dimensional subspace using singular value decomposition (SVD) Tonkin and Doherty (2005), and truncated Karhunen-Loéve expansion (KLE) Marzouk and Najm (2009) and conditional KLEs (CKLE) Li and Tartakovsky (2020). In this approach, the parameter estimation problem is reduced to estimating coefficients in the low-dimensional representation using MAP or Bayesian sampling.
@@
MCMC methods construct multiple chains of $\bm{\zeta}$ samples with stationary distributions equal to the target posterior distribution $P(\bm{\zeta}|\mathcal{D})$. However, MCMC suffers from the COD, especially the random-walk MCMC methods. One way to accelerate MCMC is to reduce autocorrelation between consecutive samples in the Markov chains. HMC Neal (2011) reduces the autocorrelation by making larger jumps in the parameter space using properties of the simulated Hamiltonian dynamics, and is a common choice for Bayesian training of statistical models Fichtner et al. (2019). However, there are challenges in applying HMC for high-dimensional IUQ problems. HMC is also not immune to COD as the increasing number of unknown parameters generally reduces the sampling efficiency of HMC and increases the autocorrelations of the consecutive samples. Another significant challenge in using HMC for IUQ problems is that the posterior covariance structure could be poorly conditioned Langmore et al. (2023) because the posterior distribution has vastly different correlations between different pairs of parameters. This can lead to biased HMC estimates Neal (2011); Betancourt (2017). Non-linearity in the model generally exacerbates the complexity of the posterior distribution and increases the HMC bias.
@@
Randomized MAP Wang et al. (2018) and similar randomization methods such as randomized maximum likelihood Chen and Oliver (2012) and randomize-then-optimize Bardsley et al. (2014) were proposed as an alternative to VI and MCMC methods for posterior sampling. For example, in randomized MAP, a (deterministic) PDE-constrained optimization problem is replaced with a stochastic problem where a stochastic objective function is minimized subject to the same deterministic PDE constraint as in the standard MAP. Then, the samples of the posterior distribution are obtained by solving the minimization problem for different realizations of the random terms in the objective function. While avoiding MCMC challenges of sampling complex posterior, the randomized MAP has the same cubic dependence on the number of unknown parameters as the standard MAP, and, therefore, is limited to the relatively low-dimensional problems.
@@
Finally, we should mention the iterative ensemble smoother (IES) and ensemble Kalman filter methods White (2018), which can be used to approximate the posterior distribution of parameters. Despite their reliance on linear regression models to approximate maps from input to output variables, these methods have demonstrated robustness even for non-linear PDE problems.
@@
In this work, we propose an efficient sampling method for high-dimensional IUQ based on the residual least-square formulation of the inverse problem found, among others, in the PICKLE and PINN methods. The main ideas of this approach include adding independent Gaussian noise to each term of the objective function in a residual least-square method and solving the resulting problem for different realizations of the noise terms. The mean and variance of the noise distributions are chosen to make the ensemble of optimization problem solutions converge to the target posterior distribution. We test this approach for the PICKLE residual least-square formulation. We apply the “randomized PICKLE” (rPICKLE) method to a 2000-dimensional IUQ problem and demonstrate that it avoids the challenges of sampling complex posterior distributions (i.e., distributions with ill-conditioned covariance matrices). Our analysis shows that the distribution of the rPICKLE samples converges to the exact posterior in the linear case with the Gaussian priors. For non-linear problems, the possible deviations from the exact posterior can be adjusted by the Metropolis algorithm. In numerical experiments, we find the deviations of the sampled distribution from the posterior distribution are minor and can be disregarded, which eliminates the need for Metropolization.
@@
This paper is organized as follows. In Section 2 we review the PICKLE formulation and present the randomized PICKLE method for approximate Bayesian parameter estimation in the PDE (1). In Section 3, we formulate rPICKLE for the IUQ problem arising in groundwater flow modeling. In Section 4, we test rPICKLE for low- (15) and high-dimensional (2000) parameter estimation problems. A comparison with HMC is provided for the low-dimensional case. Discussion and conclusions are provided in Section 5.
@@ 2.1 Conditional KLE
where $\overline{y^{c}}$ is the prior mean of $y$ conditioned on $y$ observations and $\{\lambda^{y}_{i}\}_{i=1}^{N_{\xi}}$ and $\{\psi_{i}^{y}(\mathbf{x})\}_{i=1}^{N_{\xi}}$ are the leading $N_{\xi}$ eigenvalues and eigenfunctions of the prior covariance of $y$ conditioned on $y$ observations, $C_{y}^{c}(\mathbf{x},\mathbf{x^{\prime}})$. Both $\overline{y}^{c}$ and $C_{y}^{c}(\mathbf{x},\mathbf{x^{\prime}})$ reflect the prior knowledge about $y(x)$ and are obtained via GPR equations, defined in A. In Eq. (3), $\bm{\xi}=[\xi_{1},...,\xi_{N_{\xi}}]^{T}$ is the vector of random variables with the prior independent standard normal distribution. For this prior of $\bm{\xi}$, the (prior) mean and covariance of $\hat{y}(\mathbf{x},\bm{\xi})$ is equal, up to the truncation error, to those of $y(\mathbf{x},\tilde{\omega})$. Estimating the posterior distribution of $\bm{\xi}$, i.e., the distribution constrained by the governing PDE, is the goal of rPICKLE.
@@
where $\overline{u}^{c}$ is the conditional mean of $u$, $\{\lambda_{i}^{u}\}_{i=1}^{N_{\eta}}$ and $\{\psi_{i}^{u}(\mathbf{x})\}_{i=1}^{N_{\eta}}$ are the leading eigenvalues and eigenfunctions of $C_{u}^{c}(\mathbf{x},\mathbf{x^{\prime}})$, the prior covariance of $u$ conditioned on $u$ observations. The prior covariance of $u$ is obtained by Monte Carlo sampling of the solution of Eq. (1) as described in A. As in the CKLE of $y$, $\{\eta_{i}\}_{i=1}^{N_{\eta}}$ are random variables with the independent standard normal prior distribution. The posterior distribution of $\bm{\eta}$ is estimated as part of the rPICKLE inversion.
@@ 2.3 Bayesian PICKLE
This can be achieved by taking the negative logarithm of both sides of Eq. (7) yielding
@@
$P(\mathcal{D}_{res}|\bm{\xi},\bm{\eta})=\left(\frac{1}{\sqrt{2\pi}\sigma_{r}}\right)^{N}\exp\left(-\frac{1}{2}\mathcal{R}(\mathbf{x};\bm{\xi},\bm{\eta})^{T}\Sigma_{r}^{-1}\mathcal{R}(\mathbf{x};\bm{\xi},\bm{\eta})\right).$ (12)
@@
Similarly, in Eq. (11) we can choose $C_{2}$ such that the prior is
@@
where $\bm{\omega}$, $\bm{\alpha}$, and $\bm{\beta}$ are independent random noise vectors that have the same distributions as those of $\mathcal{R}$, $\bm{\xi}$, and $\bm{\eta}$ in the Bayesian PICKLE, i.e., they have zero mean and the covariance functions $\Sigma_{w}=\Sigma_{r}={\sigma_{r}^{2}}\mathbf{I}$, $\Sigma_{\alpha}=\Sigma_{\xi}=\mathbf{I}$, and $\Sigma_{\beta}=\Sigma_{\eta}=\mathbf{I}$, respectively. Then, the samples of the posterior distribution $P(\bm{\xi},\bm{\eta}|\mathcal{D}_{res})$ are generating by solving the the minimization problem:
@@
The rPICKLE method proceeds as follows. We draw $N_{ens}$ i.i.d. samples of $\bm{\omega},\bm{\alpha},\bm{\beta}$ and, for each sample, minimize the loss (18) to obtain samples $(\bm{\xi}^{*}_{i},\bm{\eta}^{*}_{i})$ ($i=1,...,N_{ens}$), which approximate the posterior distribution of $\bm{\xi}$ and $\bm{\eta}$. Then, we use the CKLEs to obtain the samples $\hat{y}(\mathbf{x};\bm{\xi}^{*}_{i})$ and $\hat{u}(\mathbf{x};\bm{\eta}_{i}^{*})$ of the posterior distribution of $y$ and $u$, respectively. In Sections 2.4.1 and 4.1, for linear or low-dimensional problems we analytically and via a comparison with HMC show that distributions approximated with the samples $\{(\bm{\xi}^{*}_{i},\bm{\eta}^{*}_{i})\}_{i=1}^{N_{ens}}$ approach the true posteriors of $(\bm{\xi},\bm{\eta})$ with increasing $N_{ens}$ regardless of the choice of $\sigma^{2}_{r}$. In Section 4.2, we numerically demonstrate that for high-dimensional problems, rPICKLE posteriors are highly informative (we are not able to obtain HMC posterior to validate rPICKLE’s convergence to the true posterior due to the HMC’s prohibitive computational cost for the considered high-dimensional problems). We note that posterior distributions depend on the choice of $\sigma^{2}_{r}$. The value of $\sigma^{2}_{r}$ for obtaining the most informative posterior is chosen as described in Section 2.3.
@@
Next, we derive the mean and covariance of $\bm{\xi}^{*}$ and $\bm{\eta}^{*}$ given by the rPICKLE Eq. (2.4). For a linear $\mathcal{R}$, the randomized minimization problem (18) is reduced to a linear least-square problem and its solution is given by the system of linear equations:
$\frac{\partial L^{r}}{\partial\bm{\xi}}=\left[(\mathbf{A}\bm{\xi}+\mathbf{B}\bm{\eta}-\mathbf{c})^{T}\Sigma_{r}^{-1}\mathbf{A}-\bm{\omega}^{T}\Sigma_{r}^{-1}\mathbf{A}\right]+\left[(\bm{\xi}-\bm{\alpha})^{T}\Sigma_{\xi}^{-1}\right]=0$ (41)
@@
Comparing Eqs. (62) and (21) yields $\bm{\mu}=\bm{\mu}_{post}$. Next, we prove that $\Sigma=\Sigma_{post}$ is the covariance of $[\bm{\xi}^{*},\bm{\eta}^{*}]^{T}$. Denote $\bm{\zeta}=[\bm{\xi}^{*},\bm{\eta}^{*}]^{T}$. The covariance of $\bm{\zeta}$ can be calculated as
@@
2.4.2 Metropolis rejection for the non-linear model

### 23 Dec 2023 v2

@@ Abstract
We propose a physics-informed machine learning method for uncertainty quantification in high-dimensional inverse problems. In this method, the states and parameters of partial differential equations (PDEs) are approximated with truncated conditional Karhunen-Loève expansions (CKLEs), which, by construction, match the measurements of the respective variables. The maximum a posteriori (MAP) solution of the inverse problem is formulated as a minimization problem over CKLE coefficients where the loss function is the sum of the norm of PDE residuals and the $\ell_{2}$ regularization term. This MAP formulation is known as the physics-informed CKLE (PICKLE) method. Uncertainty in the inverse solution is quantified in terms of the posterior distribution of CKLE coefficients, and we sample the posterior by solving a randomized PICKLE minimization problem, formulated by adding zero-mean Gaussian perturbations in the PICKLE loss function. We call the proposed approach the randomized PICKLE (rPICKLE) method.
For linear and low-dimensional nonlinear problems (15 CKLE parameters), we show analytically and through comparison with Hamiltonian Monte Carlo (HMC) that the rPICKLE posterior converges to the true posterior given by the Bayes rule. For high-dimensional non-linear problems with 2000 CKLE parameters, we numerically demonstrate that rPICKLE posteriors are highly informative–they provide mean estimates with an accuracy comparable to the estimates given by the MAP solution and the confidence interval that mostly covers the reference solution. We are not able to obtain the HMC posterior to validate rPICKLE’s convergence to the true posterior due to the HMC’s prohibitive computational cost for the considered high-dimensional problems. Our results demonstrate the advantages of rPICKLE over HMC for approximately sampling high-dimensional posterior distributions subject to physics constraints.
@@ 1 Introduction
[…]
@@
In the inverse setting, given the measurements of $y$ and $u$, deterministic inverse solutions give the point estimates of $y$ and $u$. The standard method for finding the point estimate is to solve the PDE-constrained optimization problem where the objective function is the sum of the square differences between predicted and observed $y$ and $u$ values. These problems are ill-posed, and a unique solution cannot be found without adding regularity constraints. Adding regularization, which can be interpreted as prior knowledge, leads to the maximum a posteriori (MAP) formulation of the inverse solution. However, the PDE-constrained MAP formulations may suffer from the COD. For example, for the diffusion equation model with the unknown space-dependent diffusion coefficient considered in this work, MAP’s computational cost increases as $O(n^{3})$, where $n$ is the number of unknown parameters (discrete values of the diffusion coefficient) Yeung et al. (2022).
Models of natural and engineering systems can have thousands of (unknown) parameters, and many dimension reduction methods have been proposed for dealing with model complexity Anderson et al. (2015). For unknown parameter fields, the dimension reduction methods usually exploit the spatial correlation of the parameters. For example, in the pilot point method RamaRao et al. (1995); Doherty et al. (2010), parameters are estimated in a few preselected locations (pilot points), and the rest of the field is reconstructed using the Gaussian process regression (GPR) or “Kriging” method. Another approach is to linearly project parameters onto a low-dimensional subspace using singular value decomposition (SVD) Tonkin and Doherty (2005) and truncated Karhunen-Loéve expansion (KLE) Marzouk and Najm (2009) and conditional KLEs (CKLE) Li and Tartakovsky (2020). In this approach, the parameter estimation problem is reduced to estimating coefficients in the low-dimensional representation using MAP or Bayesian sampling.
@@
MCMC methods construct multiple chains of $\bm{\zeta}$ samples with stationary distributions equal to the target posterior distribution $P(\bm{\zeta}|\mathcal{D})$. However, MCMC suffers from the COD, especially the random-walk MCMC methods. One way to accelerate MCMC is to reduce autocorrelation between consecutive samples in the Markov chains. HMC reduces the autocorrelation by making larger jumps in the parameter space using properties of the simulated Hamiltonian dynamics and is a common choice for Bayesian training of statistical models Neal (2011); Fichtner et al. (2019). However, there are challenges in applying HMC for high-dimensional IUQ problems. HMC is also not immune to COD because the increasing number of unknown parameters generally reduces the sampling efficiency of HMC and increases the autocorrelations of the consecutive samples. Another significant challenge with using HMC for IUQ problems is that the posterior covariance structure could be poorly conditioned Langmore et al. (2023) because the posterior distribution has vastly different correlations between different pairs of parameters. This can lead to biased HMC estimates Neal (2011); Betancourt (2017). Nonlinearity in the model generally exacerbates the complexity of the posterior distribution and increases the HMC bias.
@@
Randomized MAP Wang et al. (2018) and similar randomization methods such as randomized maximum likelihood Chen and Oliver (2012) and randomize-then-optimize Bardsley et al. (2014) were proposed as alternatives to VI and MCMC methods for posterior sampling. For example, in randomized MAP, a (deterministic) PDE-constrained optimization problem is replaced with a stochastic problem where a stochastic objective function is minimized subject to the same deterministic PDE constraint as in the standard MAP. Then, the samples of the posterior distribution are obtained by solving the minimization problem for different realizations of the random terms in the objective function. While avoiding the MCMC challenges of sampling a complex posterior, the randomized MAP has the same cubic dependence on the number of unknown parameters as the standard MAP and is therefore limited to relatively low-dimensional problems.
@@
Finally, we should mention the iterative ensemble smoother (IES) and ensemble Kalman filter methods White (2018), which can be used to approximate the posterior distribution of parameters. Despite their reliance on linear regression models to approximate maps from input to output variables, these methods have demonstrated robustness even for nonlinear PDE problems.
@@
In this work, we propose an efficient sampling method for high-dimensional IUQ based on the residual least-square formulation of the inverse problem found, among others, in the PICKLE and PINN methods. The main ideas of this approach include adding independent Gaussian noise to each term of the objective function in a residual least-square method and solving the resulting problem for different realizations of the noise terms. The mean and variance of the noise distributions are chosen to make the ensemble of optimization problem solutions converge to the target posterior distribution. We test this approach for the PICKLE residual least-square formulation. We apply the “randomized PICKLE” (rPICKLE) method to a 2000-dimensional IUQ problem and demonstrate that it avoids the challenges of sampling complex posterior distributions (i.e., distributions with ill-conditioned covariance matrices). Our analysis shows that the distribution of the rPICKLE samples converges to the exact posterior in the linear case with the Gaussian priors. For nonlinear problems, the possible deviations from the exact posterior can be adjusted by the Metropolis algorithm. In numerical experiments, we find the deviations of the sampled distribution from the posterior distribution are minor and can be disregarded, which eliminates the need for Metropolization.
@@
This paper is organized as follows. In Section 2, we review the PICKLE formulation and present the randomized PICKLE method for approximate Bayesian parameter estimation in the PDE (1). In Section 3, we formulate rPICKLE for the IUQ problem arising in groundwater flow modeling. In Section 4, we test rPICKLE for low- (15) and high-dimensional (2000) parameter estimation problems. A comparison with HMC is provided for the low-dimensional case. Discussion and conclusions are provided in Section 5.
@@ 2.1 Conditional KLE
where $\overline{y^{c}}$ is the prior mean of $y$ conditioned on $y$ observations and $\{\lambda^{y}_{i}\}_{i=1}^{N_{\xi}}$ and $\{\psi_{i}^{y}(\mathbf{x})\}_{i=1}^{N_{\xi}}$ are the leading $N_{\xi}$ eigenvalues and eigenfunctions of the prior covariance of $y$ conditioned on $y$ observations, $C_{y}^{c}(\mathbf{x},\mathbf{x^{\prime}})$. Both $\overline{y}^{c}$ and $C_{y}^{c}(\mathbf{x},\mathbf{x^{\prime}})$ reflect the prior knowledge about $y(x)$ and are obtained via GPR equations, defined in A. In Eq. (3), $\bm{\xi}=[\xi_{1},...,\xi_{N_{\xi}}]^{T}$ is the vector of random variables with the prior independent standard normal distribution. For this prior of $\bm{\xi}$, the (prior) mean and covariance of $\hat{y}(\mathbf{x},\bm{\xi})$ are equal, up to the truncation error, to those of $y(\mathbf{x},\tilde{\omega})$. Estimating the posterior distribution of $\bm{\xi}$, i.e., the distribution constrained by the governing PDE, is the goal of rPICKLE.
@@
where $\overline{u}^{c}$ is the conditional mean of $u$ and $\{\lambda_{i}^{u}\}_{i=1}^{N_{\eta}}$ and $\{\psi_{i}^{u}(\mathbf{x})\}_{i=1}^{N_{\eta}}$ are the leading eigenvalues and eigenfunctions of $C_{u}^{c}(\mathbf{x},\mathbf{x^{\prime}})$, which is the prior covariance of $u$ conditioned on $u$ observations. The prior covariance of $u$ is obtained by Monte Carlo sampling of the solution of Eq. (1) as described in A. As in the CKLE of $y$, $\{\eta_{i}\}_{i=1}^{N_{\eta}}$ are random variables with the independent standard normal prior distribution. The posterior distribution of $\bm{\eta}$ is estimated as part of the rPICKLE inversion.
@@ 2.3 Bayesian PICKLE
This can be achieved by taking the negative logarithm of both sides of Eq. (7), yielding
@@
$P(\mathcal{D}_{res}|\bm{\xi},\bm{\eta})=\left(\frac{1}{\sqrt{2\pi}\sigma_{r}}\right)^{N}\exp\left(-\frac{1}{2}\mathcal{R}(\mathbf{x};\bm{\xi},\bm{\eta})^{T}\Sigma_{r}^{-1}\mathcal{R}(\mathbf{x};\bm{\xi},\bm{\eta})\right),$ (12)
@@
Similarly, in Eq. (11), we can choose $C_{2}$ such that the prior is
@@
where $\bm{\omega}$, $\bm{\alpha}$, and $\bm{\beta}$ are independent random noise vectors that have the same distributions as those of $\mathcal{R}$, $\bm{\xi}$, and $\bm{\eta}$ in the Bayesian PICKLE, i.e., they have zero mean and the covariance functions $\Sigma_{w}=\Sigma_{r}={\sigma_{r}^{2}}\mathbf{I}$, $\Sigma_{\alpha}=\Sigma_{\xi}=\mathbf{I}$, and $\Sigma_{\beta}=\Sigma_{\eta}=\mathbf{I}$, respectively. Then, the samples of the posterior distribution $P(\bm{\xi},\bm{\eta}|\mathcal{D}_{res})$ are generated by solving the the minimization problem:
@@
The rPICKLE method proceeds as follows. We draw $N_{ens}$ i.i.d. samples of $\bm{\omega}$, $\bm{\alpha}$, and $\bm{\beta}$ and, for each sample, minimize the loss (18) to obtain samples $(\bm{\xi}^{*}_{i},\bm{\eta}^{*}_{i})$ ($i=1,...,N_{ens}$), which approximate the posterior distribution of $\bm{\xi}$ and $\bm{\eta}$. Then, we use the CKLEs to obtain the samples $\hat{y}(\mathbf{x};\bm{\xi}^{*}_{i})$ and $\hat{u}(\mathbf{x};\bm{\eta}_{i}^{*})$ of the posterior distribution of $y$ and $u$, respectively. In Sections 2.4.1 and 4.1, for linear or low-dimensional problems, we analytically and via a comparison with HMC show that distributions approximated with the samples $\{(\bm{\xi}^{*}_{i},\bm{\eta}^{*}_{i})\}_{i=1}^{N_{ens}}$ approach the true posteriors of $(\bm{\xi},\bm{\eta})$ with increasing $N_{ens}$ regardless of the choice of $\sigma^{2}_{r}$. In Section 4.2, we numerically demonstrate that for high-dimensional problems, rPICKLE posteriors are highly informative (we cannot obtain the HMC posterior to validate rPICKLE’s convergence to the true posterior due to the HMC’s prohibitive computational cost for the considered high-dimensional problems). We note that posterior distributions depend on the choice of $\sigma^{2}_{r}$. The value of $\sigma^{2}_{r}$ for obtaining the most informative posterior is chosen as described in Section 2.3.
@@
Next, we derive the mean and covariance of $\bm{\xi}^{*}$ and $\bm{\eta}^{*}$ given by the rPICKLE Eq. (2.4). For a linear $\mathcal{R}$, the randomized minimization problem (18) is reduced to a linear least-square problem, and its solution is given by the system of linear equations:
$\frac{\partial L^{r}}{\partial\bm{\xi}}=\left[(\mathbf{A}\bm{\xi}+\mathbf{B}\bm{\eta}-\mathbf{c})^{T}\Sigma_{r}^{-1}\mathbf{A}-\bm{\omega}^{T}\Sigma_{r}^{-1}\mathbf{A}\right]+\left[(\bm{\xi}-\bm{\alpha})^{T}\Sigma_{\xi}^{-1}\right]=0,$ (41)
@@
Comparing Eqs. (62) and (21) yields $\bm{\mu}=\bm{\mu}_{post}$. Next, we prove that $\Sigma=\Sigma_{post}$ is the covariance of $[\bm{\xi}^{*},\bm{\eta}^{*}]^{T}$. The covariance of $\bm{\zeta}=[\bm{\xi}^{*},\bm{\eta}^{*}]^{T}$ can be calculated as
@@
2.4.2 Metropolis rejection